Two particles each of mass m are moving in a horizontal circle with same angular speed .if both string are of same length then ratio of tension in the string T1/T2=?

Grade:12

3 Answers

prabhat
6 years ago
T2 equal to Mw2(2r)And on the inner mass particle we write T1-T2=Mw2r by balance the force on T1Now put the value of T2 T1-Mw2(2r)=Mw2rT1=3Mw2r Now T1/T2=3Mw2r/2Mw2r
Yusuf
6 years ago
It`s very simple T2 is equal to centripetal force . Then T1 is equal to T2-T1=centripetal force and we can neglect velocity because it`s the same and remember for T2 r=2r and the solve you will get the ans
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

T1=[m(w*2)(r)]+[m(w*2) (2r)]
T2=[m(w*2)(2r)]

T1/T2 = 3m(w*2)(r)/2m(w*2)(r)
= 3/2

Thanks and Regards